OpenStudy (anonymous):

who knows how to do fraction patterns for 2/3, 4/5, 6/7, 8/9, ?, ?, ? i dont know how to do the last 3

OpenStudy (lgbasallote):

do you see any pattern in \[\frac 23, \; \frac 45, \; \frac 67, \; \frac 89\]

OpenStudy (anonymous):

last three are 10/11, 12/13, and 14/15

OpenStudy (anonymous):

oh im dumb

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome